OpenStudy (anonymous):

Solve: log2 (x) +log2(x+7) =3 (2 is the base)

OpenStudy (lalaly):

do you know the logarithm rules?

OpenStudy (anonymous):

Yeah i am just really bad at logarithms

OpenStudy (lalaly):

ok you know that \[ \log_a (xy) = \log_a x + \log_a y\]

OpenStudy (anonymous):

yes i do i have done that part but i dont know how to continue it from there

OpenStudy (lalaly):

ok so\[\log_2 x(x+7)=3\]

OpenStudy (lalaly):

Whenever you have\[\log_b A =x\]that means\[b^x=A\]

OpenStudy (anonymous):

yeah okay so, do you multiply the x(x-7) ?

OpenStudy (lalaly):

yep

OpenStudy (anonymous):

how about this question, 4logx(2) = log 2(x) [x and 2 are the bases] is it like this, logx(2^4) = log2 (x) log2(2^4)/ log 2(x) = log 2(x) 4/log2(x) = log2(x) how do i continue from here ?